"Ive decreased the number of calories in the food by 125% which has helped me lose weight." If the speaker used to consume 2000 calories a day, how much does he now consume? Is thia possible?
Answer by Theo(13342)

he consumed 2000 calories.

125% of 2000 calories = 2500 calories.

he now consumes 2000 - 2500 = -500 calories a day.

no, it is not possible.

if what was meant was that 2000 calories is 125% more than what he now consumes, then that would be possible.

in that case, let x = the number of calories he now consumes.

the equation becomes 1.25 * x = 2000.

solve for x to get x = 2000 / 1.25 = 1600 calories.

he is consuming 400 calories less than what he was consuming.

that would be 400 / 2500 = 16% less than what he was consuming.

16% of 2000 = .16 * 2000 = 400.

2000 - 400 = 1600.

the only way the first equation would be possible is if he stopped eating and just consumed whatever calories of energy he had in his system.

consuming -500 calories a day means he was expending 500 calories a day.

nothing was coming in.

it was all going out.